# JavisDiT [Uncaptioned image]: Joint Audio-Video Diffusion Transformer with Hierarchical Spatio-Temporal Prior Synchronization

### v1

@@ 3.3 Multi-Stage Training Strategy
Our JavisDiT focuses on two core objectives: achieving high-quality single-modal generation and ensuring fine-grained spatio-temporal alignment between generated videos and audios. To do this, we leverage the pretrained weights of OpenSora [65] for the video branch, and adopt a three-stage training strategy for robust joint video-audio generation:
@@
    Audio Pretraining. We initialize the audio branch with OpenSora’s video branch weights and train it on a large-scale audio dataset (0.8M samples) to ensure superior single-modal generation quality.
@@
    ST-Prior Training. We train the ST-Prior Estimator $\mathcal{P}_{\phi}$ using 0.6M synchronous text-video-audio triplets and synthesized asynchronous negative samples. This stage can run concurrently with the first stage.
@@
    Joint Generation. We freeze video and audio branches’ self-attention blocks and ST-Prior Estimator, training only the ST-CrossAttn and Bi-CrossAttn modules with 0.6M samples to enable synchronized video-audio generation.
In addition, dynamic temporal masking [65] is equipped to enable flexible adaptation of JavisDiT to x-conditional tasks (v2a/a2v generation, image animation, video-audio extension, etc.). Details are provided in Appendix F.
@@ 4 A Challenging JavisBench Benchmark
A strong generative model must ensure diverse video content, audio types, and fine-grained spatio-temporal synchronization. However, current JAVG benchmarks might lack diversity and robustness for comprehensive evaluation (as indicated in Sec. D.1). To combat this shortcoming, we propose a more challenging benchmark (Sec. 4.1), and also a robust metric to assess spatio-temporal alignment (Sec. 4.2).
@@ 4.1 Data Construction
Taxonomy. To assess the capabilities of joint video-audio generation models, we design five evaluation dimensions from coarse to fine:
  • Event Scenario depicts the primary scenario where the audio-visual event happens, such as nature or industry.
@@
Using GPT-4 [1], we develop a hierarchical categorization system with 5 dimensions and 19 categories (Fig. 4), and the detailed definitions are presented in Sec. D.2.
@@
Data Curation. There are two data sources: (1) test sets of existing datasets (e.g., Landscape/AIST++ and FAVDBench [46]), and (2) YouTube videos uploaded between June and December 2024 to prevent data leakage [31]. To collect YouTube videos, we prompt GPT-4 to generate category-specific keywords using our defined taxonomy, enabling efficient and targeted video collection while avoiding noisy data. After strict manual legal and ethical verification (see Appendix A), the above process yielded around 30K sounding video candidates. With several filtering tools to ensure quality and diversity, we use the advanced Qwen-family models [58, 52, 5] to generate captions and categorize video-audio pairs into desired taxonomy, and finally curate 10,140 diverse and high-quality video-audio pairs for JavisBench-10K. More details about data construction are presented in Sec. D.3. We also randomly select 1,000 samples to form the version of JavisBench-mini for efficient evaluation.
Table 1: Comparison with existing benchmarks. TAVGBench [31]’s evaluation set is not currently released (Unknown).
Dataset #Sample Scenarios Categories OpenDomain
AIST++ 20 1 (dancing) 1 (music) $\times$
Landscape 100 1 (natural) 1 (ambient) $\times$
TAVGBench 3,000 Unknown Indiscriminate $\surd$
Ours 10,140 Diversified Hierarchical $\surd$
@@
Benchmark Statistics. Tab. 1 highlights JavisBench’s contributions: (1) offering more diverse data compared to AIST++ [22] and Landscape [21], and (2) providing a detailed taxonomy for comprehensive evaluations, surpassing TAVGBench [31]. As Fig. 4 shows, JavisBench covers diverse event scenarios, visual styles, and audio types, ensuring a balanced distribution for greater diversity. We also include under-represented scenarios like 2D/3D animations (25%) and industrial events (13%) for a more realistic benchmark. 75% of samples feature multiple sounding subjects, 28% involve sequential events, and 57% include simultaneous events, presenting challenges for generative models.
Figure 4: Category distribution of our benchmark.
@@
Table 2: Main results on proposed JavisBench. Our method comprehensively outperforms or gets on par with the currently available audio/video generation models. All models generate 4-second sounding videos at 240P/24fps with 16kHz audio.
@@
FVD $\downarrow$ KVD $\downarrow$ FAD $\downarrow$ TV-IB $\uparrow$ TA-IB $\uparrow$ CLIP $\uparrow$ CLAP $\uparrow$ AV-IB $\uparrow$ CAVP $\uparrow$ AVHScore $\uparrow$ JavisScore $\uparrow$
@@
TempoTkn [60] 539.8 7.2 - 0.085 - 0.198 - 0.137 0.787 0.122 0.103
TPoS [18] 839.7 4.7 - 0.103 - 0.238 - 0.142 0.778 0.129 0.095
@@
ReWaS [19] - - 9.4 - 0.109 - 0.236 0.111 0.794 0.104 0.079
See&Hear [56] - - 7.6 - 0.072 - 0.248 0.164 0.798 0.143 0.112
FoleyC [63] - - 9.1 - 0.173 - 0.299 0.204 0.800 0.186 0.151
@@
MM-Diff [42] 2311.9 12.2 27.5 0.080 0.032 0.173 0.048 0.119 0.783 0.109 0.070
Ours 203.2 1.4 6.9 0.151 0.197 0.325 0.320 0.201 0.801 0.183 0.158
4.2 JavisScore: A More Robust JAVG Metric
Motivation. AV-Align [60] is a widely-adopted JAVG metric, which uses video optical-flow estimation to match the onset detected in audios to measure video-audio synchronization. However, AV-Align may struggle with complex scenarios (i.e., with multiple sounding events or subtle visual movements) and produce misleading results (see Sec. D.4). Therefore, we propose a more robust evaluation metric, namely JavisScore, to measure visual-audio synchronization in diverse, real-world contexts.
Implementation and Verification. Technically, we chunk each video-audio pair into several clips with 2-seconds window size and 1.5-seconds overlap, use ImageBind [11] to compute the audio-visual synchronization for each segment, and average the scores as the final metric:
where $E_{v}$ and $E_{a}$ are vision and audio encoders, $A_{i}$ and $V_{i}$ are audio and video segments in $i$-th window, and $\sigma$ is a specifically designed synchronization measure for video-audio segments. Motivated by Mao et al. [31], we calculate similarities between all frames and the audio within each segment, and select the 40% least synchronized frames to obtain the score of the current window. Sec. D.4 presents more details and discussions, where we also build a human-annotated evaluation dataset with 3,000 samples to verify the effectiveness of our JavisScore against previous metrics.
5 Experiments
5.1 Setups and Implementations
This part presents the main experimental settings, with detailed configurations provided in Sec. C.1.
Evaluation Datasets. We primarily use the proposed JavisBench as the evaluation dataset to enable comprehensive and extensive model comparisons. Additionally, we also evaluate on the AIST++ [22] and landscape [21] datasets for consistency with prior works.
Metrics. For JavisBench, we employ multiple metrics across four dimensions: (1) single-modal generation quality, (2) semantic consistency between conditional texts and generated videos/audios, (3) semantic consistency and (4) spatio-temporal synchrony between videos and audios. For the AIST++/Landscape, we follow prior works [42, 48] and report FVD, KVD, and FAD.
Compared Methods. To benchmark the audio-video generation performance on JavisBench, we reproduce and compare a series of baselines: (1) for cascaded T2A+A2V methods [60, 18], we use AudioLDM2 [27] for the prepositive T2A task; (2) for cascaded T2V+V2A methods [19, 56, 63], OpenSora [65] is adopted to generate videos at first; and (3) for JAVG models, we currently compare with MM-Diffusion [42] since others [51, 48, 26, 64] are not currently open-sourced. For AIST++ and Landscape, we directly adopt the results reported in prior works for comparison.
Figure 5: Video-audio synchrony with JavisBench’s taxonomy. Current SOTA models still suffer from challenging scenarios
Implementation Details. We collect 780K audio-text pairs from multiple sources for audio pertaining by 13 epochs, with 610K video-audio-text triplets from MMTrail [4] and TAVGBench [31] to train ST-Prior Estimator for 1 epoch and JavisDiT for 2 epochs. The learning rate is 1e-5 for ST-Prior Estimator and 1e-4 for DiT. We follow Zheng et al. [65] to leverage a bucketing strategy to support dynamic video resolution (from 144P to 1080P) and audio duration (from 2s to 16s) in batches, largely improving application flexibility during inference.
@@ 5.2 Main Results and Observations
Our JavisDiT Achieves Superior Single-Modal Quality and Video-Audio Synchrony. As shown in Tab. 2, our carefully designed spatiotemporal DiT architecture demonstrates exceptional unimodal generation quality, achieving significantly superior results compared to the UNet-based architecture (e.g., TempoToken [60]) and the naive DiT architecture (e.g., MM-Diffusion [42]), with remarkable FVD (203.2) and FAD (6.9) scores. Meanwhile, from the perspective of global semantic alignment, including text-consistency and video-audio consistency, our model also achieves state-of-the-art performance, as evidenced by its TA-IB score of 0.151 and CLIP similarity score of 0.325. Fig. 6 showcases some representative JAVG examples. Notably, in terms of audio-video synchrony, our model outperforms various cascaded and joint audio-video generation approaches, achieving a JavisScore of 0.158, surpassing the state-of-the-art cascaded method FoleyCrafter [63].
To ensure a rigorous model comparison, we follow the standard settings from prior works [42, 56, 48] and train our model for 300 epochs on two closed-set datasets, including Landscape [21] and AIST++ [22]. As demonstrated in Tab. 3, our method consistently achieves state-of-the-art performance, with FVD of 94.2 on Landscape and FAD of 9.6 on AIST++. These results further highlight the superiority of our meticulously designed DiT architecture and hierarchical spatial-temporal prior estimator.
Table 3: Experimental results on previous datasets. Numbers are borrowed from their released papers.
Method Landscape AIST++
FVD $\downarrow$ KVD $\downarrow$ FAD $\downarrow$ FVD $\downarrow$ KVD $\downarrow$ FAD $\downarrow$
MM-Diff [42] 332.1 26.6 9.9 219.6 49.1 12.3
See&Hear [56] 326.2 9.2 12.7 - - -
AV-DiT [51] 172.7 15.4 11.2 68.8 21.0 10.2
MM-LDM [48] 105.0 8.3 9.1 105.0 27.9 10.2
Ours 94.2 7.8 8.5 86.7 19.8 9.6
Current Models Fail to Simulate Complex Scenarios. Fig. 5 presents the performance of two representative methods (FoleyCrafter[63] and our JavisDiT) across JavisBench categories, showing that existing models—including ours—struggle with AV synchrony in complex scenarios. When the sounding video contains only a single sounding object (e.g., a person playing the violin alone), JavisScore is generally higher than in multi-object cases (e.g., a street performance with multiple musicians), as the latter requires identifying the correct visual-audio correspondence. Similarly, videos with multiple simultaneous events (e.g., a dog barking while a car horn sounds) yield lower JavisScore than single-event cases (e.g., a person clapping), due to the increased challenge of modeling event timing and interactions. Sec. D.5 and Fig. A7 thoroughly highlight the limitations of current JAVG in handling real-world complexity.
Figure 6: JavisDiT precisely captures the visual and auditory clues from text inputs to generate faithful sounding-videos with high-quality spatio-temporal alignments. Colored texts are spatial-temporal objects (underlined) and actions. More cases are shown in Sec. E.4.
@@ 5.3 In-depth Analyses and Discussions
To efficiently evaluate our proposed method, we perform the 3rd stage (JAVG) training using a subset of 60K entries from our entire training data, and test the models on JavisBench-mini (with 1,000 randomly selected samples from JavisBench). For simplicity, we report three normalized scores for a clear comparison from three dimensions:
  • AV-Quality: $S_{\text{AVQ}}=0.01\times S_{\text{FVD}}+S_{\text{KFD}}+0.1\times S_{\text{FAD}}$
  • AV-Consistency: $S_{\text{AVC}}=S_{\text{AV-IB}}+S_{\text{CAVP}}+S_{\text{AVHScore}}$
  • AV-Synchrony: $S_{\text{AVS}}=S_{\text{JavisScore}}$
Table 4: Ablation on module components of our JavisDiT. The specifically designed DiT architecture and ST-priors jointly contribute to single-modal quality and va-synchrony.
@@
$\times$ $\times$ $\times$ 9.371 1.140 0.133
$\surd$ $\times$ $\times$ 7.293 1.155 0.144
$\surd$ $\surd$ $\times$ 6.127 1.191 0.167
$\surd$ $\times$ $\surd$ 6.581 1.157 0.149
$\surd$ $\surd$ $\surd$ 6.012 1.201 0.169
Well-designed DiT Backbone is Superior. In Tab. 4, we first build a vanilla baseline to replace the spatio-temporal self-attention (ST-SA) backbone by UNet [27] for the audio branch, and the generation receives a pool AV-Quality at 9.371. After replacing the UNet backbone with ST-SA modules [65], $S_{\text{AVQ}}$ immediately improves to 7.293, demonstrating the efficacy of the ST-DiT architecture. Then, we respectively incorporate the spatio-temporal cross-attention (ST-CA) with our ST-Prior and the bidirectional cross-attention (Bi-CA) modules for video-audio information sharing. Accordingly, ST-CA brings significant enhancement on AV-Consistency (1.191 vs. 1.155) and AV-Synchrony (0.167 vs. 0.144), more than the simple Bi-CA module (1.157 vs. 1.155 for $S_{\text{AVC}}$ and 0.149 vs. 0.144 for $S_{\text{AVS}}$). This verifies our motivation that a simple channel-sharing mechanism in AV-DiT [51] cannot effectively build audio-video synchronization, which can be achieved by our proposed fine-grained ST-Prior guidance instead. After bridging ST-SA, ST-CA, and Bi-CA modules, our JavisDiT reaches the best performance for both single-modal quality ($S_{\text{AVQ}}=6.012$) and video-audio synchronization ($S_{\text{AVC}}=1.201$ and $S_{\text{AVS}}=0.169$), demonstrating the superiority of the well-designed DiT backbone.
Spatio-Temporal Prior is Effective and Generic. In Tab. 5, we take a preliminary step to investigate the number of spatio-temporal priors and the way to utilize ST-Priors for better video-audio synchronization. We first take the model without ST-CA modules (the 4th row in Tab. 4) as our baseline (the 1st row in Tab. 5), and add ST-CA modules by gradually increasing the prior number from 1 to 32. According to the results, both single-modal quality ($S_{\text{AVQ}}$) and video-audio synchrony ($S_{\text{AVC}}$ and $S_{\text{AVS}}$) consistently increase as the prior number gains. Then, we try to utilize the 32 spatial-temporal priors for addition (adding to video/audio latent representations like a conditional embedding) and modulation (mapping priors to scales and biases to modulate video/audio representations). Although the final performance is inferior to the utilization of cross-attention, ST-priors still considerably enhance all metrics (e.g., 0.159/0.161 vs. 0.149 for $S_{\text{AVS}}$). Besides, we further experiment and evaluate the influence of prior numbers and dimensions in Fig. A8, and verify the optimization objectives during ST-priors’ estimation in Tab. A6. All the empirical results verify the effectiveness and versatility of our estimated spatio-temporal priors.
Table 5: Ablation on token number and utilization of ST-priors. We provide more experimental results in Fig. A8 and Tab. A6.
@@
0 0 - 6.581 1.157 0.149
1 1 CrossAttn 6.909 1.188 0.163
16 16 CrossAttn 6.322 1.200 0.166
32 32 CrossAttn 6.012 1.201 0.169
32 32 Addition 6.267 1.183 0.159
32 32 Modulate 6.190 1.191 0.161

### v2

@@ 3.3 Multi-Stage Training Strategy
Our JavisDiT focuses on two core objectives: achieving high-quality single-modal generation and ensuring fine-grained spatio-temporal alignment between generated videos and audios. To do this, we leverage the pretrained weights of OpenSora (Zheng et al., 2024) for the video branch, and adopt a three-stage training strategy for robust joint video-audio generation:
@@
    Audio Pretraining. We initialize the audio branch with OpenSora’s video branch weights and train it on 0.8M audio-text samples to ensure superior single-modal generation quality.
@@
    ST-Prior Training. We train the ST-Prior Estimator $\mathcal{P}_{\phi}$ using 0.6M synchronous text-video-audio triplets and synthesized asynchronous negative samples to strengthen representation.
@@
    JAVG Training. We freeze the video and audio branches’ self-attention blocks and ST-Prior Estimator, training only the ST-CrossAttn and Bi-CrossAttn modules with 0.6M samples to enable synchronized video-audio generation.
In addition, dynamic temporal masking (Zheng et al., 2024) enables flexible adaptation of JavisDiT to x-conditional tasks (v2a/a2v generation, image animation, video-audio extension, etc.). Details refer to Sec. E.7.
@@ 4 A Challenging JavisBench Benchmark
A strong generative model must ensure diverse video content, audio types, and fine-grained spatio-temporal synchrony. However, current JAVG evaluation benchmarks fall short in encompassing the diversity of scenes, the complexity of real-world environments, and the presence of multiple sounding events within a single audio-video instance, thereby constraining comprehensive and realistic evaluation (detailed in Sec. D.1). To combat this, we propose a more challenging benchmark featuring complex multi-event video-audio pairs (Sec. 4.1), along with a robust synchronization-oriented metric tailored for evaluating fine-grained spatio-temporal alignment (Sec. 4.2).
@@ 4.1 Data Construction
Taxonomy. We design five evaluation dimensions from coarse to fine:
  • Event Scenario depicts scenarios where audio-visual events happen, such as nature or industry.
@@
Using GPT-4 (Achiam et al., 2023), we develop a hierarchical categorization system with 5 dimensions and 19 categories (Fig. 4), and the detailed definitions are presented in Sec. D.2.
@@
Data Curation. There are two data sources: (1) test sets of existing datasets (e.g., Landscape/AIST++ and FAVDBench (Shen et al., 2023)), and (2) YouTube videos uploaded between June and December 2024 to prevent data leakage (Mao et al., 2024). To collect YouTube videos, we prompt GPT-4 to generate category-specific keywords using our defined taxonomy, enabling efficient and targeted video collection while avoiding noisy data. After strict manual legal and ethical verification (see Sec. A.1), the above process yielded around 30K sounding video candidates. With several filtering tools to ensure quality and diversity, we use the advanced Qwen-family models (Yang et al., 2024a; Wang et al., 2024c; Chu et al., 2024) to generate captions and categorize video-audio pairs into desired taxonomy, and finally curate 10,140 diverse and high-quality video-audio pairs for JavisBench-10K. More details about data construction are presented in Sec. D.3. We also randomly select 1,000 samples to form the version of JavisBench-mini for efficient evaluation.
Figure 4: Left: Comparison between JavisBench and other benchmarks. TAVGBench (Mao et al., 2024)’s test set is not yet released (Unknown). Right: Detailed category distribution of JavisBench.
@@
Benchmark Statistics. Fig. 4 highlights JavisBench’s contributions: (1) offering more diverse data compared to AIST++ (Li et al., 2021) and Landscape (Lee et al., 2022), (2) providing a detailed taxonomy for comprehensive evaluations, surpassing TAVGBench (Mao et al., 2024), and (3) featuring the first evaluation benchmark focused on multi-event synchronization. As Fig. 4 shows, JavisBench covers diverse event scenarios, visual styles, and audio types, ensuring a balanced distribution for greater diversity. Most realistic but under-represented scenarios like 2D/3D animations (25%) and industrial events (13%) are included. Furthermore, JavisBench emphasizes spatial and temporal complexity. 75% of samples feature multiple sounding events, 28% involve sequential events, and 57% include simultaneous events. These diverse multi-event scenarios pose significant challenges for JAVG models, particularly in achieving accurate spatial and temporal alignment.
4.2 JavisScore: A More Robust JAVG Metric
Motivation. AV-Align (Yariv et al., 2024) is a widely-adopted JAVG metric, which uses video optical-flow estimation to match the onset detected in audios to measure temporal synchronization. However, AV-Align may struggle with complex scenarios (i.e., with multiple sounding events or subtle visual movements) and produce misleading results (see Sec. D.4). Therefore, we propose a more robust evaluation metric, namely JavisScore, to measure spatiotemporal synchronization in diverse, real-world contexts.
Implementation and Verification. Technically, we chunk each video-audio pair into several clips with 2-second window size and 1.5-second overlap, use ImageBind (Girdhar et al., 2023) to compute the audio-visual synchronization for each segment, and average the scores as the final metric:
where $E_{v}$ and $E_{a}$ are vision and audio encoders, $A_{i}$ and $V_{i}$ are audio and video segments in $i$-th window, $W$ is the number of windows, and $\sigma$ is a specifically designed synchronization measure for video-audio segments (see Sec. D.4). Motivated by Mao et al. (2024), we calculate similarities between all frames and the audio within each segment, and select the 40% least synchronized frames to obtain the score of the current window. Sec. D.4 also presents a human-annotated evaluation dataset with 3,000 samples to verify JavisScore’s efficacy against previous metrics.
5 Experiments
5.1 Setup
Evaluation Datasets and Metrics. The proposed JavisBench is used as the primary benchmark for comprehensive evaluation. Multiple metrics are reported: (1) audio/video quality, (2) semantic consistency against conditional texts, (3) semantic consistency and (4) spatio-temporal synchrony between videos and audios. We also evaluate on AIST++ (Li et al., 2021) and landscape (Lee et al., 2022) for consistency, and follow prior works (Ruan et al., 2023; Sun et al., 2024) to report FVD, KVD, and FAD.
Compared Methods. For JavisBench, we reproduce and compare a series of baselines: (1) for cascaded T2A+A2V methods (Yariv et al., 2024; Jeong et al., 2023), we use AudioLDM2 (Liu et al., 2024b) for the prepositive T2A task; (2) for cascaded T2V+V2A methods (Jeong et al., 2024; Xing et al., 2024; Zhang et al., 2024), OpenSora (Zheng et al., 2024) is adopted to generate videos at first; and (3) for JAVG models, we currently compare with MM-Diffusion (Ruan et al., 2023) and UniVerse-1 Wang et al. (2025) since the others (Liu et al., 2024a; Zhao et al., 2025) are not currently open-sourced. For AIST++ and Landscape, we directly adopt the reported results.
Implementation Details. We collect 780K audio-text pairs for audio pertaining by 55 epochs, with 610K video-audio-text triplets to train ST-Prior Estimator for 1 epoch and JavisDiT for 2 epochs. The learning rate is 1e-5 for ST-Prior Estimator and 1e-4 for DiT. The video encoder-decoder and audio encoder-decoder are taken from OpenSora (Zheng et al., 2024) and AudioLDM2 (Liu et al., 2024b) and kept frozen. Detailed configurations refer to Sec. C.1.
@@ 5.1 Setup
Table 1: Main results on proposed JavisBench. Models are evaluated with 4-second videos at 240P/24fps and audio at 16kHz. JavisDiT comprehensively outperforms or gets on par with currently available SoTA approaches. Best and secondary results are marked bold and underline respectively.
@@
[…]
@@
TempoTkn (Yariv et al., 2024) 539.8 7.2 - 0.084 - 0.205 - 0.137 0.787 0.122 0.103
TPoS (Jeong et al., 2023) 839.7 4.7 - 0.201 - 0.229 - 0.142 0.778 0.129 0.095
@@
ReWaS (Jeong et al., 2024) - - 9.4 - 0.123 - 0.280 0.110 0.794 0.104 0.079
See&Hear (Xing et al., 2024) - - 7.6 - 0.129 - 0.263 0.160 0.798 0.143 0.112
FoleyCftr (Zhang et al., 2024) - - 9.1 - 0.149 - 0.383 0.193 0.800 0.186 0.151
@@
MM-Diff (Ruan et al., 2023) 2311.9 12.2 27.5 0.080 0.014 0.181 0.079 0.119 0.783 0.109 0.070
UniVerse-1 (Wang et al., 2025) 194.2 0.5 8.7 0.272 0.111 0.309 0.245 0.104 0.793 0.098 0.077
JavisDiT(Ours) 204.1 1.4 7.2 0.263 0.143 0.302 0.391 0.197 0.801 0.179 0.154
Table 2: Experimental results on previous datasets. Numbers are borrowed from their released papers, and our model still performs best.
Method Landscape AIST++
[…]
$M$-Diff 332.1 26.6 9.9 219.6 49.1 12.3
See&Hear 326.2 9.2 12.7 - - -
AV-DiT 172.7 15.4 11.2 68.8 21.0 10.2
$M$-LDM 105.0 8.3 9.1 105.0 27.9 10.2
JavisDiT(Ours) 94.2 7.8 8.5 86.7 19.8 9.6
Figure 5: Generative av-synchrony with JavisBench’s taxonomy. Current SOTA models still suffer from challenging scenarios.
@@ 5.2 Main Results and Observations
Our JavisDiT achieves superior single-Modal quality and video-audio synchrony. As shown in Tab. 1, our carefully designed spatiotemporal DiT architecture demonstrates exceptional unimodal generation quality, achieving significantly superior results compared to UNet-based architectures (e.g., TempoToken (Yariv et al., 2024)) and naive DiT architectures (e.g., MM-Diffusion (Ruan et al., 2023)), with remarkable FVD (204.1) and FAD (7.2) scores. Meanwhile, from the perspective of global semantic alignment, including text-consistency and video-audio consistency, our model also achieves state-of-the-art performance, as evidenced by its TA-IB score of 0.263 and CLIP similarity score of 0.302. We attribute the gap between our JavisDiT and UniVerse-1 (Wang et al., 2025) in video-text alignment to the gap of pretrained models (we use OpenSora while UniVerse-1 takes more powerful Wan2.1 (Wan et al., 2025)). Fig. 6 showcases some representative JAVG examples. Notably, in terms of audio-video synchrony, our end-to-end model outperforms various cascaded and joint audio-video generation approaches, achieving a JavisScore of 0.154, surpassing the state-of-the-art cascaded method FoleyCrafter (Zhang et al., 2024).
To ensure a rigorous comparison, we follow the standard settings from prior works (Ruan et al., 2023; Xing et al., 2024; Sun et al., 2024) and train our model for 300 epochs on two closed-set datasets, including Landscape (Lee et al., 2022) and AIST++ (Li et al., 2021). As demonstrated in Tab. 2, our method consistently achieves state-of-the-art performance, with FVD of 94.2 on Landscape and FAD of 9.6 on AIST++. These results further highlight the superiority of our meticulously designed DiT architecture and hierarchical spatial-temporal prior estimator.
Current models fail to simulate complex scenarios. Tab. 2 presents FoleyCrafter(Zhang et al., 2024) and our JavisDiT across JavisBench categories, showing that existing models — including ours — struggle with AV synchrony in complex scenarios. When the sounding video contains only a single sounding object (e.g., a person playing the violin alone), JavisScore is generally higher than in multi-object cases (e.g., a street performance with multiple instruments), as the latter requires identifying the correct visual-audio correspondence. Similarly, videos with multiple simultaneous events (e.g., a dog barking while a car horn sounds) yield lower JavisScore than single-event cases (e.g., a person clapping), due to the increased challenge of modeling event timing and interactions. Sec. D.5 and Fig. A9 thoroughly highlight the limitations in handling real-world complexity.
@@ 5.3 In-depth Analyses and Discussions
To efficiently evaluate our proposed method, we perform the 3rd stage (JAVG) training using a subset of 60K entries from our entire training data, and test the models on JavisBench-mini (with 1,000 randomly selected samples from JavisBench). For simplicity, we report three normalized scores for a clear comparison from three dimensions: (1) Quality: $S_{\text{AVQ}}=0.01\times S_{\text{FVD}}+S_{\text{KFD}}+0.1\times S_{\text{FAD}}$; (2) Consistency: $S_{\text{AVC}}=S_{\text{AV-IB}}+S_{\text{CAVP}}+S_{\text{AVHScore}}$; (3) Synchrony: $S_{\text{AVS}}=S_{\text{JavisScore}}$.
Figure 6: JavisDiT precisely captures the visual and auditory clues from text inputs to generate faithful sounding-videos with high-quality spatio-temporal alignments. Colored texts are spatio-temporal objects (underlined) and actions. More cases are shown in Sec. E.6.
Table 3: Ablation on the model design. The specifically developed DiT architecture and the HiST-Sypo estimator jointly contribute to single-modal quality and va-synchrony. Full evaluation results are presented in Tab. A5.
@@
$\times$ $\times$ $\times$ 9.371 1.140 0.118
$\surd$ $\times$ $\times$ 7.293 1.155 0.130
$\surd$ $\surd$ $\times$ 6.127 1.191 0.150
$\surd$ $\times$ $\surd$ 6.581 1.157 0.133
$\surd$ $\surd$ $\surd$ 6.012 1.201 0.153
Table 4: Ablation on token number and injection strategies of ST-Priors. We provide more experimental results in Fig. A10 and Tab. A8. Full evaluation results are presented in Tab. A6.
@@
0 0 - 6.581 1.157 0.133
1 1 CrossAttn 6.909 1.188 0.137
16 16 CrossAttn 6.322 1.200 0.151
32 32 CrossAttn 6.012 1.201 0.153
32 32 Addition 6.267 1.183 0.144
32 32 Modulate 6.190 1.191 0.145
Well-designed DiT backbone is superior. In Tab. 4, we first build a vanilla baseline to replace the spatio-temporal transformer backbone by UNet (Liu et al., 2024b) for the audio branch, and the generation receives a pool AV-Quality at 9.371. After replacing the UNet backbone with STDiT modules (Zheng et al., 2024), $S_{\text{AVQ}}$ immediately improves to 7.293, demonstrating the efficacy of the STDiT architecture. Then, we respectively incorporate the HiST-Sypo estimator and the bidirectional cross-attention (BiCA) modules for video-audio information sharing. Accordingly, HiST-Sypo brings significant enhancement on AV-Consistency (1.191 vs. 1.155) and AV-Synchrony (0.150 vs. 0.130), more than the simple BiCA module (1.157 vs. 1.155 for $S_{\text{AVC}}$ and 0.133 vs. 0.130 for $S_{\text{AVS}}$). This verifies our motivation that a simple channel-sharing mechanism in AV-DiT (Wang et al., 2024b) cannot effectively build audio-video synchronization, which can be achieved by our proposed fine-grained ST-Prior guidance instead. After bridging STDiT, HiST-Sypo, and BiCA modules, our JavisDiT reaches the best performance for both single-modal quality ($S_{\text{AVQ}}=6.012$) and video-audio synchronization ($S_{\text{AVC}}=1.201$ and $S_{\text{AVS}}=0.153$), demonstrating the superiority of the well-designed DiT backbone.
Spatio-temporal prior is effective and generic. In Tab. 4, we take a preliminary step to investigate the number of spatio-temporal priors and the way to inject ST-Priors for better video-audio synchronization. We first take the model without ST-CA modules (the 4th row in Tab. 4) as our baseline (the 1st row in Tab. 4), and add ST-CA modules by gradually increasing the prior number from 1 to 32. According to the results, both single-modal quality ($S_{\text{AVQ}}$) and video-audio synchrony ($S_{\text{AVC}}$ and $S_{\text{AVS}}$) consistently increase as the prior number gains. Then, we try to utilize the 32 spatial-temporal priors for addition (adding to video/audio latent representations like a conditional embedding) and modulation (mapping priors to scales and biases to modulate video/audio representations). Although the final performance is inferior to the utilization of cross-attention, ST-priors still considerably enhance all metrics (e.g., 0.144/0.145 vs. 0.133 for $S_{\text{AVS}}$). Besides, we further experiment and evaluate the influence of prior numbers and dimensions in Fig. A10, and verify the optimization objectives during ST-priors’ estimation in Tab. A8. All the empirical results verify the effectiveness and versatility of our estimated spatio-temporal priors.